How do I do this problem?
One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. I nned to find the larger number.
Thank you, Jamie

Answer by stanbon(48545)

One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. I nned to find the larger number.
-----------
Let one of the numbers be "x".
The larger number is "x+12".
EQUATION:
x+2(x+12)=39
x+2x+24=39
3x=15
x=5 (this is the smaller number)
x+12=17 (this is the larger number)
Cheers,
Stan H.